OpenStudy (anonymous):

Evaluate 2X TO THE 2 POWER + 1 FOR X = 3? SHOW WORK

OpenStudy (angela210793):

2x^2+1 2*3^2+1=2*9+1=18+1=19

OpenStudy (anonymous):

is this right

OpenStudy (anonymous):

yeah that is right.